Solve -dx/x(x+y) = dy/y(x+y) = dz/(x-y)(2x+2y+z)
1
Expert's answer
2020-11-02T20:40:26-0500

We have "\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)} = \frac{dz}{(x-y)(2x+2y+z)}" ______________(1)

Now from first two part of equation (1),

"\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)}"

"\implies \frac{dx}{-x} = \frac{dy}{y}"

Now by integration, we get "-\ln(x) = \ln(y) - \ln(c_1) \implies c_1 = xy" ________(2)

Now also "\frac{dx}{-x(x+y)} = \frac{dy}{y(x+y)} = \frac{dz}{(x-y)(2x+2y+z)} = \frac{2dx+2dy+dz}{(x-y)z}"

"\implies \frac{dz}{(x-y)(2x+2y+z)} = \frac{2dx+2dy+dz}{(x-y)z}"

"\implies \frac{dz}{2x+2y+z} = \frac{2dx+2dy+dz}{z}"

"\implies zdz = (2x+2y+z) (2dx+2dy+dz)"

Hence by integration, we get

"\frac{z^2}{2} = \frac{(2x+y+z)^2}{2} + \frac{c_2}{2}"

"\implies c_2 = z^2- (2x+y+z)^2" ______________(3)

Hence, solution is :

"c_2 = f(c_1)"

"\implies z^2- (2x+y+z)^2 = f(xy)" .